The pH of a 0.14 M solution of an unknown weak acid is 5.85. Calculate the Ka of the acid.
Answer #1

Concentration=0-14M PH_ 5.85 pH = + (pka_logc) 5.85 = 1 (pka log (OM) 11-1= pka_log (6.14) pka = 11:17 log (0.14) = 10.85 = 1